Based on a past record, a supermarket finds that 52% of people who enter the store will make a purchase. 24 people enter the store in an hour. (a) Find the mean and variance of customers who make a purchase. (b) Give ONE real life example of the probability is very unlikely to occur. (c) Give ONE real life example of the probability is very likely to occur. 


1
Expert's answer
2021-03-16T07:10:13-0400

Here,

Probability of people who make purchase p=52%=0.52

Probability of people who do not make purchase q=1-p=1-0.52=0.48


Number of people enter the store in an hour n = 24


(a) Mean = "np=24\times 0.52=12.48"

Variance = "npq=24\times 0.52\times 0.48=5.9904"


(b) Rates of car accidents have increased rapidly in the past decades. For example, if a city has a population of one lakh, and the death rate in car accidents is 500. So, the chance of being killed in a crash is 500/1 lakh is 0.05%. Thus, a person has a 0.05% chance to die in a car accident.


(C) Flipping a coin is one of the most important events before the start of the match. There is no surety, either head will come or not. Both head and tail have 1 out of 2, i.e., 50% chances to occur.